# Defensive Model Expansion for Robust Bayesian Inference

### v1

@@ 3.3 Results for Gaussian Processes
Theorem 7 shows that a certain “information-weighted” projection $\beta^{\star}$ satisfies a semiparametric Bernstein–von Mises theorem with $I$ the efficient Fisher information under $\mathcal{F}_{0}$. The quantity $\widehat{\beta}$ is equivalent to the MLE under the parametric logistic regression submodel, being a one-step Newton–Raphson update of a $\sqrt{N}$-consistent quantity. We use the information-weighted projection mainly for technical convenience, since it makes $\beta^{\star}$ a linear functional of $r(\cdot)$. In practice, it is much better to use the Kullback-Leibler projection $\beta^{\star}=\arg\min_{\beta}\sum_{i}\mu(X_{i})\log\mu(X_{i})/\mu_{\beta}(X_{i})$ where $\mu_{\beta}(x)=(1+e^{-x^{\top}\beta})^{-1}$, as in experiments we have found it to be far less biased.
@@ Results
Figure 2 shows the results. We see that when $\lambda_{0}=0$, the performance of GBART is indistinguishable from that of the linear model at all sample sizes. Somewhat surprisingly, for smaller signal-to-noise ratios and sample sizes the unmodified BART model outperforms a correctly-specified linear model; this is because the sample size is not large enough to estimate the coefficients reliably, giving an advantage to models that regularize the predictions towards zero. This defect in the linear model can be mitigated by replacing the least squares estimates with a ridge regression or lasso.
When $\lambda_{0}=0.4$, GBART outperforms the other methods, particularly at large sample sizes. It is strictly superior to the linear model under all data generating mechanisms. We again see that when both the sample size and the signal-to-noise ratios are small there is an advantage to using a BART model that more strongly regularizes the predictions.
@@
Figure 3 shows that the spike-and-GP prior behaves as expected. When $\lambda_{0}=0$ the parametric model is consistently selected, while as $\lambda_{0}$ increases the nonparametric model is selected with high probability. The trend is more muted when $\sigma_{0}$ is larger; this is intuitive, since (for example) $\sigma_{0}=4$ corresponds to a signal-to-noise ratio of roughly 16-to-1, while $\sigma_{0}=1$ corresponds to a signal-to-noise ratio of roughly 1-to-1.
@@
4.3 Semiparametric Bernstein–von Mises for Gaussian Processes
We now empirically validate the Bernstein–von Mises result for the projection of the GP posterior onto the linear regression model. We use the data generating process
@@ 4.3 Semiparametric Bernstein–von Mises for Gaussian Processes
with $X_{ij}\sim\operatorname{Normal}(0,1)$ and $\epsilon_{i}\sim\operatorname{Normal}(0,1)$ independently. We set $P=5$ and $\beta_{0j}\equiv 1.55$. We then fit a GP regression model $Y_{i}=\mu(X_{i})+\epsilon_{i}$ with $\mu\sim\operatorname{GP}(0,\kappa)$ and compute the posterior distribution of the projection parameter $\beta^{\star}=(\boldsymbol{X}^{\top}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top}\boldsymbol{\mu}$. We consider the Laplace kernel $\kappa(x,x^{\prime})=e^{-\|x-x^{\prime}\|}$, the squared exponential kernel $\kappa(x,x^{\prime})=e^{-\|x-x^{\prime}\|^{2}}$, and the squared-exponential-plus-linear kernel $\kappa(x,x^{\prime})=100\,(x^{\prime})^{\top}x+e^{-\|x-x^{\prime}\|_{2}}$.
@@
We monitor three quantities: (i) the nominal coverage of 95% credible intervals for $\beta^{\star}_{1}$ under each kernel; (ii) the scaled bias of the posterior mean, defined as the average of $\sqrt{N}(\widehat{\beta}_{1}-\beta_{01})$ with $\widehat{\beta}=\mathbb{E}_{\Pi}(\beta^{\star}\mid\boldsymbol{Z}_{N})$; and (iii) the scaled variance, defined as the average of $N\operatorname{Var}_{\Pi}(\beta^{\star}_{1}\mid\boldsymbol{Z}_{N})$. If the posterior satisfies a semiparametric Bernstein–von Mises result, we expect coverage near 95%, scaled bias near zero, and scaled variance near 1, since $\mathbb{E}_{\theta_{0}}(\boldsymbol{X}^{\top}\boldsymbol{X})=\mathrm{I}$ in this setup.
@@ Results
Figure 4 shows results across kernels and sample sizes. Among the three, only the squared-exponential-plus-linear kernel empirically satisfies the semiparametric Bernstein–von Mises theorem at the chosen sample sizes. In larger samples, we do see the asymptotic variance of the squared exponential and Laplace kernels approaching the semiparametric efficiency, and indeed asymptotically we do attain the semiparametric efficiency level; the problem with these kernels is that, even in very large samples, the estimates are highly biased, which is clearly seen in the middle panel of Figure 4.
@@ 4.4 Diagnosing Model Misspecification in Practice
We now show how to diagnose model misspecification in practice by analyzing data from the Medical Expenditure Panel Survey (MEPS, Cohen et al.,, 2009). We consider a subset of the data from the 2011 survey containing 16,113 with the goal of predicting self-assessed health status. We use Model NR with $Y_{i}$ representing the self-assessed health status of individual $i$ and $X_{i}$ representing the individual’s age, body mass index (BMI), education level, income relative to the poverty line, region of the country, sex, marital status, race, seatbelt use, and smoking status.
To illustrate the ability of the GBART model to default to a linear model when there is little data while also adapting to non-linearities and interactions in the data, we consider two analyses: one using a randomly selected sample of size of $N=1000$ and one using a sample of size $N=8000$. For both cases we monitor the summary $R^{2}$ of the projection onto a linear regression. To assess how closely the semiparametric Bernstein-von Mises result holds, we also monitor the posterior distribution of the coefficient of BMI.
@@
Results for the $N=1000$ model are given in Figure 5. We see that the model is shrunk very heavily towards a linear regression model, with the posterior summary $R^{2}$ being very close to $100\%$. The posterior distribution of the coefficient of BMI is also very close to the posterior from a parametric Bayesian linear model, indicating that the semiparametric Bernstein-von Mises approximation is accurate. Finally, the projected linear regression model gives predictions that are virtually indistinguishable from the full general BART model. This indicates that, at $N=1000$, there is limited evidence of model misspecification so that the general BART model defaults to the linear model.
@@
From the analysis for the $N=8000$ setting, we see that there is evidence of misspecification of the linear regression, implying that there exist either non-linearities or interactions in the data. An advantage of the fully-Bayesian approach is that we are free to summarize the posterior in any way we choose, and in particular we can look for better models to project the posterior of $\mu(x)$ on to. To this end, in Figure 7 we give the summary $R^{2}$ and projected predictions for a model that includes spline terms for age, BMI, poverty level, with interaction terms for (age) $\times$ BMI and (age) $\times$ (poverty level). After making this modification, we find that the overall summary $R^{2}$ of the model exceeds 98%, with the posterior distribution of $R^{2}$ centered around 96%, giving a roughly $10\%$ increase in predictive variability accounted for.
Figure 8: Decision tree fit to the residuals $\widehat{\mu}(X_{0})-\widetilde{\mu}(X_{i})$ from the additive model.
To gain better insight into the remaining sources of variability, Figure 8 displays a CART fit to the residuals $\widehat{\mu}(X_{i})-\widetilde{\mu}(X_{i})$, where $\widetilde{\mu}(x)$ denotes the additive summary of $\widehat{\mu}(x)$. The magnitude of the predictions of the CART model represent how off, on average, the additive summary is from the general BART model for subsets of the observations. The largest degree of misfit occurs for individuals who are white and with poverty level less than $121$, with the additive summary predicting a health status on average 0.036 units higher than the GBART model.
@@ 5 Discussion
The position advocated for in this manuscript is one of defensive model expansion, wherein we embed a parametric Bayesian model inside a nonparametric Bayesian model and apply strong shrinkage towards the parametric model. We claim that this retains many of the benefits of specifying a parametric Bayesian model, such as strong performance when $N$ is small or the signal-to-noise ratio is low, while mitigating the drawbacks of potential model misspecification.

### v2

@@ 3.3 Results for Gaussian Processes and Posterior Projections
Theorem 7 shows that a certain “information-weighted” projection $\beta^{\star}$ satisfies a semiparametric Bernstein-von Mises theorem with $I$ the efficient Fisher information under $\mathcal{F}_{0}$. The quantity $\widehat{\beta}$ is asymptotically equivalent to the MLE under the parametric logistic regression submodel, being a one-step Newton–Raphson update of a $\sqrt{N}$-consistent quantity. We use the information-weighted projection mainly for technical convenience, as it makes $\beta^{\star}$ a linear functional of $\theta(x)=r(x)+x^{\top}\beta$. In practice, it is much better to use the Kullback-Leibler projection $\beta^{\star}=\arg\min_{\beta}\sum_{i}\mu(X_{i})\log\frac{\mu(X_{i})}{\mu_{\beta}(X_{i})}+\{1-\mu(X_{i})\}\log\frac{1-\mu(X_{i})}{1-\mu_{\beta}(X_{i})}$ where $\mu_{\beta}(x)=(1+e^{-x^{\top}\beta})^{-1}$, as in experiments we have found it to be far less biased.
@@ Results
Figure 2 shows the results. We see that when $\lambda_{0}=0$, the performance of GBART is indistinguishable from that of the linear model at all sample sizes. Somewhat surprisingly, for smaller signal-to-noise ratios and sample sizes the unmodified BART model outperforms a correctly-specified linear model; this is because the sample size is not large enough to estimate the coefficients reliably, giving an advantage to models that regularize the predictions toward zero. This defect in the linear model can be mitigated by replacing the least squares estimates with a ridge regression or lasso.
When $\lambda_{0}=0.4$, GBART outperforms the other methods, particularly at large sample sizes. Under these simulation settings, it is strictly superior to the linear model under all data generating mechanisms. We again see that when both the sample size and the signal-to-noise ratios are small there is an advantage to using a BART model that more strongly regularizes the predictions.
@@
Figure 3 shows that the spike-and-GP prior behaves as expected. When $\lambda_{0}=0$ the parametric model is consistently selected, while as $\lambda_{0}$ increases the nonparametric model is selected with high probability. The trend is more muted when $\sigma_{0}$ is larger; this is intuitive, since (for example) $\sigma_{0}=4$ corresponds to a signal-to-noise ratio of roughly 1-to-16, while $\sigma_{0}=1$ corresponds to a signal-to-noise ratio of roughly 1-to-1.
@@
4.3 Semiparametric Bernstein-von Mises for Gaussian Processes
We now empirically validate the Bernstein-von Mises result for the projection of the GP posterior onto the linear regression model. We use the data generating process
@@ 4.3 Semiparametric Bernstein-von Mises for Gaussian Processes
with $X_{ij}\sim\operatorname{Normal}(0,1)$ and $\epsilon_{i}\sim\operatorname{Normal}(0,1)$ independently. We set $P=5$ and $\beta_{0j}\equiv 1.55$. We then fit a GP regression model $Y_{i}=\mu(X_{i})+\epsilon_{i}$ with $\mu\sim\operatorname{GP}(0,\kappa)$ and compute the posterior distribution of the projection parameter $\beta^{\star}=(\boldsymbol{X}^{\top}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top}\boldsymbol{\mu}$. We consider the Laplace kernel $\kappa(x,x^{\prime})=e^{-\|x-x^{\prime}\|}$, the squared exponential kernel $\kappa(x,x^{\prime})=e^{-\|x-x^{\prime}\|^{2}}$, and the squared-exponential-plus-linear kernel $\kappa(x,x^{\prime})=100\,(x^{\prime})^{\top}x+e^{-\|x-x^{\prime}\|_{2}^{2}}$.
@@
We monitor three quantities: (i) the nominal coverage of 95% credible intervals for $\beta^{\star}_{1}$ under each kernel; (ii) the scaled bias of the posterior mean, defined as the average of $\sqrt{N}(\widehat{\beta}_{1}-\beta_{01})$ with $\widehat{\beta}=\mathbb{E}_{\Pi}(\beta^{\star}\mid\boldsymbol{Z}_{N})$; and (iii) the scaled variance, defined as the average of $N\operatorname{Var}_{\Pi}(\beta^{\star}_{1}\mid\boldsymbol{Z}_{N})$. If the posterior satisfies a semiparametric Bernstein-von Mises result, we expect coverage near 95%, scaled bias near zero, and scaled variance near 1, since $\mathbb{E}_{\theta_{0}}(N^{-1}\boldsymbol{X}^{\top}\boldsymbol{X})=\mathrm{I}$ in this setup.
@@ Results
Figure 4 shows results across kernels and sample sizes. Among the three, only the squared-exponential-plus-linear kernel empirically satisfies the semiparametric Bernstein-von Mises theorem at the chosen sample sizes. In larger samples, we do see the asymptotic variance of the squared exponential and Laplace kernels approaching the semiparametric efficiency, and indeed asymptotically we do attain the semiparametric efficiency level; the problem with these kernels is that, even in very large samples, the estimates are highly biased, which is clearly seen in the middle panel of Figure 4.
@@ 4.4 Diagnosing Model Misspecification in Practice
We now show how to diagnose model misspecification in practice by analyzing data from the Medical Expenditure Panel Survey (MEPS, Cohen et al.,, 2009). We consider a subset of the 2011 survey with 16,113 observations, with the goal of predicting self-assessed health status. We use Model NR with $Y_{i}$ representing the self-assessed health status of individual $i$ and $X_{i}$ representing the individual’s age, body mass index (BMI), education level, income relative to the poverty line, region of the country, sex, marital status, race, seatbelt use, and smoking status. Although self-assessed health status is ordinal, here we use Model NR as a simple working Gaussian model for prediction and projection-based summarization; extending the same workflow to ordinal or logistic-style models is also possible.
To illustrate the ability of the GBART model to default to a linear model when there is little data while also adapting to nonlinearities and interactions in the data, we consider two analyses: one using a randomly selected sample of size $N=1000$ and one using a sample of size $N=8000$. For both cases we monitor the summary $R^{2}$ of the projection onto a linear regression. All projection summaries reported below are in-sample summaries computed on the analyzed sample. To assess how closely the semiparametric Bernstein-von Mises result holds, we also monitor the posterior distribution of the coefficient of BMI.
@@
Results for the $N=1000$ model are given in Figure 5. We see that the model is shrunk very heavily toward a linear regression model, with the posterior summary $R^{2}$ being very close to $100\%$. The posterior distribution of the coefficient of BMI is also very close to the posterior from a parametric Bayesian linear model, indicating that the semiparametric Bernstein-von Mises approximation is accurate. Finally, the projected linear regression model gives predictions that are virtually indistinguishable from the full general BART model. This indicates that, at $N=1000$, there is limited evidence of model misspecification so that the general BART model defaults to the linear model.
@@
From the analysis for the $N=8000$ setting, we see that there is evidence of misspecification of the linear regression, implying that there are either nonlinearities or interactions in the data. An advantage of the fully-Bayesian approach is that we are free to summarize the posterior in any way we choose, and in particular we can look for better models to project the posterior of $\mu(x)$ onto. To this end, in Figure 7 we give the summary $R^{2}$ and projected predictions for a model that includes spline terms for age, BMI, poverty level, with interaction terms for (age) $\times$ BMI and (age) $\times$ (poverty level). After making this modification, we find that the overall summary $R^{2}$ of the model exceeds 98%, with the posterior distribution of $R^{2}$ centered around 96%, giving a roughly $10\%$ increase in predictive variability accounted for.
Figure 8: Decision tree fit to the residuals $\widehat{\mu}(X_{i})-\widetilde{\mu}(X_{i})$ from the additive model.
To gain better insight into the remaining sources of variability, Figure 8 displays a CART fit to the residuals $\widehat{\mu}(X_{i})-\widetilde{\mu}(X_{i})$, where $\widetilde{\mu}(x)$ denotes the additive summary of $\widehat{\mu}(x)$. The CART predictions quantify the average discrepancy between the additive summary and the GBART fit within each subgroup. The largest degree of misfit occurs for white individuals with poverty level below $121$, with the additive summary predicting a health status on average 0.036 units higher than the GBART model.
@@ 5 Discussion
The position advocated for in this manuscript is one of defensive model expansion, wherein we embed a parametric Bayesian model inside a nonparametric Bayesian model and apply strong shrinkage toward the parametric model. We claim that this retains many of the benefits of specifying a parametric Bayesian model, such as strong performance when $N$ is small or the signal-to-noise ratio is low, while mitigating the drawbacks of potential model misspecification.